OpenStudy (anonymous):

if Tan(theta)=x, then Sin(2*theta) is equal to: the possible answers are: a) 2x b) 1+x^2 c) (1+x^2)/2 d) (2x)/(1+x^2) Please help

OpenStudy (anonymous):

possible answer "c" is wrong. The correct answer is "(1+x^2)/2x"

OpenStudy (anonymous):

you need to use \[\sin(2x)=2\sin(x)\cos(x)\]

OpenStudy (anonymous):

Ok let me try

OpenStudy (anonymous):

you need of course to find \[\sin(x)\] and \[\cos(x)\]

OpenStudy (anonymous):

I got a really weird thing. This is what I did. sen(2theta)=2sin(theta)cos(theta) Tan(theta)=sen(thetha)/Cos(theta) sen(theta)=xCos(theta) sen(2theta)=2XCos(theta)Cos(theta) sen(2theta)=2XCos2(theta)

OpenStudy (anonymous):

Im really bad at this. can you show the work

OpenStudy (anonymous):

draw a triangle and convince yourself that if \[tan(\theta)=x\] then \[\sin(\theta)=\frac{x}{\sqrt{1+x^2}}\] and \[\cos(\theta)=\frac{1}{\sqrt{1+x^2}}\]

OpenStudy (anonymous):

ahhh by pythagoras. Clear.

OpenStudy (anonymous):

Should I Apply sen(2thta)=2Sin(theta)Cos(theta)? I think so

OpenStudy (anonymous):

Satellite can you help me with the question I just posted, please?

OpenStudy (anonymous):

This question in already solved. I need help in my lattest question

OpenStudy (anonymous):

ya you can apply 2 sin thta * cos thta.............but it takes too much time....just apply sin(2tha)= (2tan tha)/(1 + tan^2 thta)... if you apply this than your answer becomes (d)

myininaya (myininaya):

the answer is d as in deliciousness